Complete Job-Order Costing EYK 3-1

EYK3-1. Business Decision Case Elizabeth Flanigan and Associates is an engineering and design firm that specializes in developing plans for recycling plants for municipalities. The firm uses a job costing system to accumulate the cost associated with each design project. Flanigan employs three levels of employee: senior engineers, associate engineers, and clerical staff. The salary cost of the senior engineers and the associate engineers is assigned to each project as direct labor. The salary cost of the clerical staff is included in overhead, along with the cost of engineering supplies, automobile travel, and equipment depreciation. The cost of airline travel, motels, building permits, and fees from other consultants is charged to each project as direct materials. Overhead is applied to projects using a predetermined overhead rate based on total engineering hours. The rate for 20X0 is $5 per hour.

The six different salary levels for 20X1 for the employees of Elizabeth Flanigan and Associates are listed below. The hourly rate is determined by dividing the yearly salary by 2,000 hours per year.

Senior engineer

Level 1: $44,000 per year ($22 per hour)

Level 2: $36,000 per year ($18 per hour)

Associate engineer

Level 3: $30,000 per year ($15 per hour)

Level 4: 24,000 per year ($12 per hour)

Clerical Staff

Level 5: $16,000 per year ($8 per hour)

     Level 6: $14,500 per year ($7.25 per hour)

The billings that are sent to the municipalities for engineering services utilize cost-plus billing. Typically, the total costs accumulated for a project (direct materials, direct labor, and overhead) are multiplied by 140% to determine the amount of the billing. The differences between the billed amount and the accumulated cost is the “plus” in cost plus.

During March 20X1, Flanigan accumulated the following information related to Job 295 for Johnson Creek City:

Senior engineer hours

Level 1……………………………………………………….52

Level 2……………………………………………………..84

Associate engineer hours

Level 3………………………………………………………106

Level 4………………………………………………………..44

Clerical hours

Level 5…………………………………………………………20

Level 6……………………………………………………..…66

Building Permits……………………………………$1,500

Airline travel and motel………………….……….$865

Required

  1. What amount should be billed to Johnson Creek City for March 20X1?
  2. How much profit was earned on Job 295 during March 20X1

PLEASE ANSWER QUESTIONS IN THIS FORM:

A) Direct Materials:    B)

Building Permits

Airline Travel and Motel

Direct Labor:

senior engineers

Overhead

Total Project Cost

Cost Plus Percentage

Total Amount billed
Answer #1
DIRECT MATERIAL
BUILDING PERMIT $    1,500.00
AIRLINE TRAVEL AND MOTEL $        865.00
DIRECT LABOR
SENIOR ENGINEERS
LEVEL 1 52*22 $    1,144.00
LEVEL 2 84*18 $    1,512.00
ASSOCIATE ENGINEER HOURS
LEVEL 3 106*15 $    1,590.00
LEVEL 4 44*12 $        528.00
OVERHEAD $    1,430.00
286 * $ 5.00
TOTAL COST $    8,569.00
BILLING AMOUNT 140% OF TOTAL COSTS $ 11,996.60
PROFITS $    3,427.60

SHOW FORMULA IN EXCEL:

DIRECT MATERIAL
BUILDING PERMIT 1500
AIRLINE TRAVEL AND MOTEL 865
DIRECT LABOR
SENIOR ENGINEERS
LEVEL 1 52*22 =52*22
LEVEL 2 84*18 =84*18
ASSOCIATE ENGINEER HOURS
LEVEL 3 106*15 =106*15
LEVEL 4 44*12 =44*12
OVERHEAD =286*5
286 * $ 5.00
TOTAL COST =SUM(B2:B15)
BILLING AMOUNT 140% OF TOTAL COSTS =+B17*1.4
PROFITS =+B18-B17